OpenStudy (cutiecomittee123):

Krypton-91 is a radioactive substance that decays very quickly. The function Q(t)=Qoe^-kt models radioactive decay of krypton-91. Q represents the quantity remaining after t seconds and the decay constant k is approximately 0.07. How long will it take a quantity of krypton-91 to decay t o10% of its origional amount? round your answer to the nearest second.

OpenStudy (anonymous):

Cant tell if this is chemistry or math

OpenStudy (cutiecomittee123):

Its math! precalc:)

OpenStudy (anonymous):

Well can yah help with this?

OpenStudy (anonymous):

i giv medal fan and testimony

OpenStudy (cutiecomittee123):

dude I cant I dont have time sorry :(

OpenStudy (anonymous):

awrh

OpenStudy (perl):

Lets suppose the original amount Qo is 100 grams of the radioactive substance

OpenStudy (cutiecomittee123):

Okay

OpenStudy (perl):

Lets suppose the original amount Qo is 100 grams of the radioactive substance. Therefore the amount of substance left at time t is modeled by the equation: Q(t) = 100*e^(-.07*t) We want to find how long will it take to decay to 10% of its original amount. so what is 10% of 100 grams? that is 10 grams So we need to solve this equation: 10 = 100 * e^(-.07 * t )

OpenStudy (cutiecomittee123):

alright so first thing I would do is divide both sides by 100:)

OpenStudy (perl):

right

OpenStudy (cutiecomittee123):

I get -1.43 seconds after dividing by 100 on both sides I get .1=e^-0.07t then divide by -0.07 on both sides and so .1/-0.07=-1.43 when rounded up. am I right??

OpenStudy (cutiecomittee123):

@perl

OpenStudy (perl):

I agree with your work up to here .1=e^-0.07t

OpenStudy (perl):

now you can't divide both sides by -.07, since -.07 is an exponent

OpenStudy (cutiecomittee123):

oh then I have to take ln(.1)/-0.07

OpenStudy (perl):

right $$ \Large .1 = e^ {-.07t} \\~\\ $$

OpenStudy (perl):

t = ln (.1) / -.07

OpenStudy (cutiecomittee123):

Okay so after doing that I get t= -32.89

OpenStudy (perl):

t should be positive

OpenStudy (cutiecomittee123):

but we cant have negative seconds

OpenStudy (perl):

ln(.1) / (-.07) = 32.89

OpenStudy (cutiecomittee123):

ah I redid it and got that! thanks:)) I forgot to do -.07